OpenStudy (perl):

help< i need help with a 3d grapher, i am using www.romanlabs.com I cannot figure out the code how to graph simple curves like f(x,y)= x^2 + y^2

OpenStudy (anonymous):

can you see rotations using this program

OpenStudy (perl):

im not sure, i cant graph even the most basic of basic functions

OpenStudy (anonymous):

is it free

OpenStudy (perl):

yes of course

OpenStudy (perl):

its a 30 day trial ,

OpenStudy (perl):

it just looks like it has a sweet interface and what not , its a shame i cant figure it out

OpenStudy (anonymous):

looks good, i will try it

OpenStudy (anonymous):

where do you enter the equations

OpenStudy (perl):

go to graph in the menu, and press ADD function graph

OpenStudy (anonymous):

then how do you enter the functions

OpenStudy (perl):

thats as far as i got, lol

OpenStudy (perl):

what a strange program ,

OpenStudy (anonymous):

man it look o cool, but i guess i need to read the instructions

OpenStudy (perl):

where are they?

OpenStudy (anonymous):

i cant find them

OpenStudy (anonymous):

it would be cool to see rotations about the x and y axis

OpenStudy (perl):

yeah, doesnt seem to rotate but you can type y = sin u

OpenStudy (anonymous):

but it should

OpenStudy (perl):

oh i get it, its like parametric

OpenStudy (perl):

ok so let x = u, y = sin (u)

OpenStudy (anonymous):

but where do i type in y=sin(u)

OpenStudy (perl):

press Ctrl+F

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then where do i go

OpenStudy (perl):

where it says x(u,v,t) = u y(u,v,t) = sin u

OpenStudy (perl):

use parenthesees

OpenStudy (perl):

sin(u)

OpenStudy (anonymous):

hey that is too cool, try rotating the graph

OpenStudy (perl):

ok i found it

OpenStudy (perl):

F6

OpenStudy (anonymous):

let me see if i can get like a circle or cone

OpenStudy (anonymous):

have you found out how to enter that equation, you posted

OpenStudy (perl):

yes now i can x(u,v,t) = u y(u,v,t) = v z(u,v,t) = u^2 + v^2

OpenStudy (anonymous):

yup

OpenStudy (perl):

ok, , thats cool

OpenStudy (anonymous):

yes that one is

OpenStudy (anonymous):

what math are you working on?

OpenStudy (perl):

calc of two variables

OpenStudy (anonymous):

man, but i dont think it generates 3D rotation

OpenStudy (perl):

it does there should be a clockwise arrow in the pane above

OpenStudy (anonymous):

yeah but if i enter like x^2 or something, i wanted to see like a 3D anmination of the cone it hould produce when rotated about the y axis. But its still easier to imagine it using this program

OpenStudy (anonymous):

maybe if i plot points

OpenStudy (perl):

i dont think it can rotate like what youre thinking

OpenStudy (perl):

like upside down revolving

OpenStudy (anonymous):

how can i create those 3d graphs from the graph gallary page

OpenStudy (perl):

well you can create the z = u^2 + v^2

OpenStudy (anonymous):

i wanna create these grpahs: http://www.romanlab.com/3dg/graphs.htm

OpenStudy (perl):

yeah, its like you have to figure it out from scratch

OpenStudy (anonymous):

i will play around with this program and see what i do, but for the momement it seem really similar to graphmatica

OpenStudy (perl):

whats that

OpenStudy (anonymous):

its a similar program: http://graphmatica.com/